OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (15, 35) and (195, 50), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points?

OpenStudy (anonymous):

Okay. So you know how ordered pairs are set up, right? (x, y) So, for example, take a look at the FIRST ordered pair: (15, 35) In this case... X = 15 Y = 35 You see? ~~~~~~~~~~~~~ Now, we have two ordered pairs. Here's the formula for finding the slope: \[\frac{ y_{1}-y_{2} }{x_{1}-x_{2} }\] ~~~~~~~~~~~~~~~ I'll walk you through this formula step by step. 1. Look at our ordered pairs. (15, 35) and (195, 50) 2. Find the FIRST Y, and subtract the SECOND Y. 35 - 50 = -15 This is the NUMERATOR (top) of the slope fraction. 3. Find the FIRST X, and subtract the SECOND X. 15 - 195 = -180 This is the DENOMINATOR (bottom) of the slop fraction. ~~~~~~~~~~~~~~ So now we have the slope fraction! \[\frac{ -15 }{ -180 }\] which is the same as \[\frac{ 15 }{ 180 }\]

OpenStudy (anonymous):

Thanks!